OpenStudy (anonymous):

how do you get 12 and 5 when you factor 3x^2 + 7x - 20? could someone show me how to solve it? thank you

OpenStudy (anonymous):

there is a couple ways to do this

OpenStudy (anonymous):

one way is to plug it into the quadratic equation

jimthompson5910 (jim_thompson5910):

Multiply 3 and -20 to get -60 find two numbers that multiply to -60 and add to 7. These two numbers are 12 and -5

OpenStudy (anonymous):

do you know the quadratic equation

hero (hero):

3x^2 + 7x - 20 = 3x^2 + 12x - 5x - 20 = 3x(x + 4) - 5(x + 4) = (x + 4)(3x - 5) You get 12x and -5x by splitting the middle term 7x since 12 - 5 = 7